# Redefining Machine Unlearning: A Conformal Prediction-Motivated Approach

### v1

@@ 4.2 Measure Machine Unlearning Method via New Metrics
RT $0.1$ $14.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $98.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $86.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $0.775({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.786({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $19.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.6})$ $98.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $86.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $0.729({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.163})$ $0.786({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.114})$ $26.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}11.7})$ $98.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.1})$ $85.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.2})$ $0.649({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.243})$ $0.765({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.135})$
$0.2$ $0.934({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.935({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.898({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.108})$ $0.932({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.133})$ $0.839({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.049})$ $0.929({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.130})$
FT $0.1$ $6.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}7.8})$ $97.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.9})$ $84.1\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.9})$ $0.791({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.087})$ $0.685({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.201})$ $9.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.9})$ $97.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.4})$ $83.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.4})$ $0.753({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.139})$ $0.683({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.217})$ $13.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.9})$ $97.2\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.6})$ $83.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.4})$ $0.718({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.174})$ $0.683({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.217})$
$0.2$ $0.969({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.905({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.053})$ $0.942({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.152})$ $0.893({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.094})$ $0.914({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.124})$ $0.890({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.091})$
RL $0.1$ $26.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}12.2})$ $96.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.8})$ $81.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.6})$ $0.338({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.540})$ $0.488({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.398})$ $31.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}17.1})$ $95.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}17.9})$ $80.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.1})$ $0.278({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.614})$ $0.451({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.449})$ $36.2\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}21.5})$ $95.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.5})$ $80.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.6})$ $0.254({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.638})$ $0.449({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.451})$
$0.2$ $0.821({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.143})$ $0.848({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.110})$ $0.752({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.038})$ $0.825({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.026})$ $0.718({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.072})$ $0.827({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.028})$
MIACR Metric.
In Table 6, we show the MIACR results on CIFAR-10 under both $10\%$ and $50\%$ random data forgetting scenarios. Compared to MIACR, traditional MIA remains a privacy concern. Albeit MIA cannot predict some forget data points as training data, it still has high confidence in appearing in the prediction set. For example, under evaluation criterion ➊, the teacher’s MIA is the method closest to the RT method, but it performs poorly in the CR metric. Similarly, FF and Salun are the top two unlearning methods in the traditional MIA metric (ideally small) under evaluation criterion ➋, but FF has low MIACR (ideally greater) values in both 10% and 50% forgetting scenarios.
4.3 Performance of Our Unlearning Framework
In this experiment, we apply the RT, FT, and RL methods to our framework. Table 7 presents the results for CIFAR-10 with ResNet-18 and Tiny ImageNet with ViT in $10\%$ random data forgetting scenario.
@@ 4.3 Performance of Our Unlearning Framework
From the perspective of evaluation criterion ➊, the FT method achieves an average gap of $1.47$ across UA, RA, and TA on ResNet-18 when $\lambda=0.5$, compared to an average gap of $2.2$ when $\lambda=0$. Similarly, on the ViT model, FT reduces the average gap from $3.53$ to $1.63$ when $\lambda=0.5$. If UA of an unlearning method is already lower than that of RT, e.g., RL, our framework cannot further improve their performance under evaluation criterion ➊.
@@ 5 Related Work
Machine unlearning has emerged as a vital research topic due to several privacy, regulatory, and ethical concerns associated with machine learning models. It refers to the process of selectively removing specific data points from a trained machine-learning model. Generally, post-hoc machine unlearning can be divided into training-based (Graves et al., 2021; Warnecke et al., 2021; Thudi et al., 2022; Tarun et al., 2023) and training-free approaches (Guo et al., 2019; Sekhari et al., 2021; Nguyen et al., 2020; Golatkar et al., 2020b, 2021; Foster et al., 2024).
To evaluate these methods, several unlearning metrics have been proposed, including UA (Brophy & Lowd, 2021; Foster et al., 2024) and MIA (Shokri et al., 2017; Chen et al., 2021). However, these metrics often fail to account for the confidence of the forgetting quality. To address this limitation, we improve it in our work motivated by conformal prediction (Angelopoulos & Bates, 2021), which stands out among uncertainty quantification techniques for its ability to provide well-calibrated, reliable confidence measures. As a generic methodology, conformal prediction can transform the outputs of any black box prediction algorithm into a prediction set. Due to its versatility, many works have specifically designed numerous conformal prediction methods tailored to particular prediction problems (Papadopoulos et al., 2002; Lei & Wasserman, 2014; Romano et al., 2020; Lei et al., 2018).
One work (Becker & Liebig, 2022) has primarily focused on parameter-level uncertainty without fully addressing the broader implications of unlearning on prediction confidence. It assesses the sensitivity of model parameters to the target data through Fisher Information Matrix, but they often rely on computationally intensive operations and may struggle to scale to large models or datasets.
@@ 6 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Setting Details
For CIFAR-10 with a ResNet-18 architecture, we train the original model from scratch for 200 epochs using SGD with a Cosine Annealing learning rate schedule, starting from an initial learning rate of 0.1. We set momentum to 0.9 and a batch size of 64. For the ViT architecture, we initialize the original model by training a pretrained ViT model for 15 epochs on Tiny ImageNet. We start with a learning rate of 0.001, while other training parameters match those used for ResNet-18. In class-wise forgetting, in each dataset, we designate the same class as the forget data $\mathcal{D}_{f}$ across all unlearning methods, with the remaining classes used as the retain data $\mathcal{D}_{r}$.
@@ B.1 Mis-label Number and In-set Ratios
Conformal prediction is applied to UA and MIA predictions to determine the number of misclassified data points (mis-label) and the number of these points that fall within the conformal prediction set (in-set) across 9 unlearning methods. We evaluate both the UA and MIA metrics by counting the misclassified data points and calculating how many of them are included in the conformal prediction set. The detailed results are presented in Table 8.
B.2 CR Metric
Tables 9 and 10 show the unlearning performance of seven unlearning methods on CIFAR-10 with ResNet-18 in $10\%$ and $50\%$ random data forgetting scenarios, while Table 11 is the results in class-wise forgetting scenario. Tables 12 and 13 present the unlearning performance on Tiny ImageNet with ResNet-18 in random data forgetting scenario, while Table 11 details the unlearning performance in class-wise forgetting scenario.
For all machine unlearning methods, increasing the $\alpha$ level leads to lower Coverage and smaller Set Size. Conversely, the CR tends to grow as $\alpha$ increases. Most methods yield a Set Size of less than $1$ when $\alpha$ is set too low, since conformal prediction is fundamentally linked to the model’s baseline prediction performance (accuracy). If the model’s accuracy significantly exceeds the confidence level, conformal prediction can easily achieve the required coverage. In fact, it may produce empty prediction sets for certain data points while still satisfying the coverage target. Excessively high $\alpha$ values can distort evaluation results, preventing the CR metric from accurately reflecting the model’s performance. Therefore, the choice of $\alpha$ should be carefully considered in relation to the model’s inherent performance. The more experimental results for random data forgetting scenarios have already been analyzed in detail in Section 4 and will not be reiterated here.

### v2

@@ 4.3 Performance of Our Unlearning Framework
CPU-RT $14.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $98.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $86.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $0.503({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.516({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $19.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.6})$ $98.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $86.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $0.458({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.045})$ $0.516({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $26.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}11.7})$ $98.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.1})$ $85.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.2})$ $0.396({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.107})$ $0.489({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.027})$
CPU-FT $6.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}7.8})$ $97.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.9})$ $84.1\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.9})$ $0.466({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.037})$ $0.389({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.127})$ $9.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.9})$ $97.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.4})$ $83.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.4})$ $0.441({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.062})$ $0.399({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.117})$ $13.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.9})$ $97.2\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.6})$ $83.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.4})$ $0.413({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.090})$ $0.401({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.115})$
CPU-RL $26.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}12.2})$ $96.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.8})$ $81.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.6})$ $0.054({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.449})$ $0.111({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.405})$ $31.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}17.1})$ $95.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}17.9})$ $80.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.1})$ $0.051({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.452})$ $0.111({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.405})$ $36.2\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}21.5})$ $95.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.5})$ $80.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.6})$ $0.051({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.452})$ $0.121({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.395})$
(a) $\mathcal{D}_{f}$
(b) $\mathcal{D}_{r}$
(c) $\mathcal{D}_{test}$
(d) $\mathcal{D}_{f}$
(e) $\mathcal{D}_{r}$
(f) $\mathcal{D}_{test}$
Figure 3: CPU-FT accuracy of $\mathcal{D}_{f}$, $\mathcal{D}_{r}$ and $\mathcal{D}_{test}$ under different $\lambda$ values across each epoch on CIFAR-10 (a-c) and Tiny ImageNet (d-f). As $\lambda$ increases, accuracy on $\mathcal{D}_{f}$ drops significantly, while retain and test accuracy remain stable.
@@
Moreover, in Figure 3, we further present the CPU-FT accuracy on forget data $\mathcal{D}_{f}$, retain data $\mathcal{D}_{r}$ and test data $\mathcal{D}_{test}$ under different $\lambda$ values across each epoch on Tiny ImageNet with ViT for $10\%$ random data forgetting. As $\lambda$ increases, the accuracy on $\mathcal{D}_{f}$ drops quickly, showing stronger unlearning effectiveness, while the accuracy on $\mathcal{D}_{r}$ and $\mathcal{D}_{test}$ remains stable. In summary, the experimental results demonstrate that our framework notably enhances the forgetting quality while maintaining stable predictive performance.
@@ 5 Related Work
Machine unlearning has emerged as a vital research topic due to several privacy, regulatory, and ethical concerns associated with machine learning models. It refers to the process of selectively removing specific data points from a trained machine learning model. Generally, post-hoc machine unlearning can be divided into training-based Graves et al. (2021); Tarun et al. (2023); Thudi et al. (2022); Warnecke et al. (2021) and training-free approaches Foster et al. (2024); Golatkar et al. (2021, 2020); Guo et al. (2019); Nguyen et al. (2020); Sekhari et al. (2021).
To evaluate these methods, several unlearning metrics have been proposed, including UA Brophy and Lowd (2021); Foster et al. (2024) and MIA Shokri et al. (2017); Chen et al. (2021). However, these metrics often fail to account for the confidence of the forgetting quality. To address this limitation, we improve it in our work motivated by conformal prediction Angelopoulos and Bates (2021), which stands out among uncertainty quantification techniques for its ability to provide well-calibrated, reliable confidence measures. As a generic methodology, conformal prediction can transform the outputs of any black box prediction algorithm into a prediction set. Due to its versatility, many works have specifically designed numerous conformal prediction methods tailored to particular prediction problems Lei et al. (2018); Lei and Wasserman (2014); Papadopoulos et al. (2002); Romano et al. (2020).
One work Becker and Liebig (2022) has primarily focused on parameter-level uncertainty without fully addressing the broader implications of unlearning on prediction confidence. It assesses the sensitivity of model parameters to the target data through Fisher Information Matrix, but they often rely on computationally intensive operations and may struggle to scale to large models or datasets.
@@ 6 Conclusion
Appendix
@@ Appendix A Setting Details
For CIFAR-10 with ResNet-18 architecture, we train the original model from scratch for 200 epochs using SGD with a Cosine Annealing learning rate schedule, starting from an initial learning rate of 0.1. We set the momentum to 0.9 and a batch size of 64. The RT model adopts the same training configuration. Other models are trained for the following durations: FT for 20 epochs, RL for 10 epochs, SalUn for 10 epochs, GA for 1 epoch (to avoid over-forgetting and significant RA degradation), NegGrad+ for 10 epochs (reduced to 2 epochs in class-wise scenarios), and SFRon for 10 epochs. All other hyperparameters match those of the original model.
For the ViT architecture, we initialize the original model by training a pretrained ViT model for 15 epochs on Tiny ImageNet. We start with a learning rate of 0.001, while other training parameters match those used for ResNet-18. We use SGD and set the momentum to 0.9 and a batch size of 64. The RT model follows the same training procedure as the original model. Other models are trained for the following durations: FT for 5 epochs, RL for 5 epochs, SalUn for 5 epochs, GA for 1 epoch, NegGrad+ for 5 epochs, and SFRon for 5 epochs. All other hyperparameters are consistent with the original model’s training.
For CIFAR-10/Tiny ImageNet, we randomly select $200$/$50$ data points per class ($2000$/$10000$ data points in total) as calibration data $\mathcal{D}_{c}$ and $\mathcal{D}_{c}^{\prime}$, respectively. The calibration data $\mathcal{D}_{c}$ does not participate in the model training or unlearning processes and is only used for calibrating the threshold $\hat{q}$, while $\mathcal{D}_{c}^{\prime}$ is used in the process of our unlearning framework to generate $\bar{q}$.
All experiments are conducted on 1 Tesla V100-SXM2 GPU card with 32GB memory in a single node.
@@ B.1 Mis-label Number and In-set Ratios
SFRon 423 121 28.6% 1,871 433 23.1%
Conformal prediction is applied to UA and MIA predictions to determine the number of misclassified data points (mis-label) and the number of these points that fall within the conformal prediction set (in-set) across 9 unlearning methods. We evaluate both the UA and MIA metrics by counting the misclassified data points and calculating how many of them are included in the conformal prediction set. The detailed results are presented in Table 6, which is the extended results of Table 2.
B.2 Distribution Comparison of Forgotten Data on UA and CR
As shown in Figures 7-10, we further analyze the probability and loss distributions of ground truth labels for data identified as truly forgotten by CR (i.e., out-set) and UA (i.e., mis-label) respectively. The distribution curves are fitted using KDE for clearer visualization. The softmax outputs for ‘out-set’ are consistently near 00 compared to ‘mis-label’, which strongly suggests that ‘out-set’ more rigorously captures real forgotten data. In cross entropy loss distribution, forgotten data identified by CR consistently show higher cross entropy loss than UA. Higher loss indicates better forgetting quality, which further validates that CR better remove fake unlearning data.
B.3 CR Metric
Tables 7 and 8 show the unlearning performance of 9 unlearning methods on CIFAR-10 with ResNet-18 in $10\%$ and $50\%$ random data forgetting scenarios, while Table 9 is the results in class-wise forgetting scenario. Tables 10 and 11 present the unlearning performance on Tiny ImageNet with ResNet-18 in random data forgetting scenario, while Table 12 details the unlearning performance in class-wise forgetting scenario. For class-wise forgetting scenario, we note $\mathcal{D}_{test}=\mathcal{D}_{tf}\cup\mathcal{D}_{tr}$. $\mathcal{D}_{tf}$ corresponds to the test-forget data exclusively containing the forget class, while $\mathcal{D}_{tr}$ represents the test-retain data within the test data $\mathcal{D}_{test}$.
For all unlearning methods, as $\alpha$ level increases, it results in reduced Coverage and smaller Set Size. This happens because a higher $\alpha$ loosens the conformal threshold $\hat{q}$, allowing fewer predictions to be included within the prediction set for each data point. On the contrary, the CR tends to increase with increasing $\alpha$. Although both Coverage and Set Size may decrease, Set Size often decreases more significantly. Consequently, the CR value of $\mathcal{D}_{f}$ generally becomes larger as $\alpha$ increases. It is natural that the adjustment of $\alpha$ affects both Coverage and Set Size. However, the final CR value really depends on the model’s performance itself. For a strict evaluation, we encourage setting $\alpha$ to $0.5$.
When $\alpha$ is set to $0.2$, most methods show a value of Set Size less than $1$ in both Table 7, 8, 10, 11. The intuition behind it is that conformal prediction, as a static predictor, is intrinsically tied to the model’s base prediction performance and accuracy. When the model’s accuracy is significantly higher than the confidence level, conformal prediction can achieve the required coverage with ease. In fact, it can generate partial empty prediction sets for some data points while still meeting the target coverage. Thus, the choice of $\alpha$ is crucial. Overly high $\alpha$ values may skew evaluation results by failing to let CR accurately reflect model performance. Therefore, we emphasize that a small $\alpha$ is generally appropriate for most unlearning scenarios.